OpenStudy (anonymous):

Hi guys, one question with a couple of parts! What is the oxidation state for the atoms in the reaction: 2Al + 8H2O → 2H+1 + 3H2 + 2Al(OH)4-1 a) Elemental aluminum (Al) b) One hydrogen atom in H2O c) One oxygen atom in H20 d) Elemental hydrogen (H) e) One aluminum atom in Al(OH)4-1 f) One oxygen atom in Al(OH)4-1 g) One hydrogen atom in Al(OH)4-1

OpenStudy (aaronq):

oxidation state on homonuclear molecules or elemental species = 0

OpenStudy (aaronq):

Oxygen = -2 always

OpenStudy (aaronq):

you can figure out the rest by balancing the charges

OpenStudy (anonymous):

Alright, so for a i got 0, b -3, c -2, d 1, e 2, f -2 and g 0. Are these right?

OpenStudy (frostbite):

Let me see if I get this right: You suggest that hydrogen in water have the oxidation number -3?

OpenStudy (anonymous):

I believe so

OpenStudy (frostbite):

Okay. We need an additional rule: "The sum of oxidations numbers most equal to total charge of the molecule." So when we look at water it have no charge... sense oxygen always get -2 each hydrogen most be what?

OpenStudy (anonymous):

-5? And in that equation, which element was reduced and which was oxidized?

OpenStudy (anonymous):

And can verify my answers please?

OpenStudy (frostbite):

Best you get some training, so I'm gonna start with water and we take them each by each. H2O we take notice: Total charge = 0 oxygen always -2 Hydrogen always +1 This also make sense we should say: 0=-2+2*x <-> x=+1

OpenStudy (frostbite):

aluminum is right. again it does not have charge and is an elemental species.

OpenStudy (anonymous):

Both hydrogens are positive one as well?

OpenStudy (frostbite):

Yes. it is 2 times +1

OpenStudy (anonymous):

Alright, so what about for e through g?

OpenStudy (anonymous):

hello?

OpenStudy (frostbite):

Sorry I was thinking. both e and g are wrong.

OpenStudy (frostbite):

The oxidation state of oxygen is -2 the oxidation state of hydrogen is +1: so we get: Al(OH)4 -1 -1 = -8+4+x Here is x the oxidation state of Al.. What is x?

OpenStudy (frostbite):

Not quite.

OpenStudy (frostbite):

\[-1=-8+4+x\] \[x=-1+8-4=3\]

OpenStudy (anonymous):

So aluminum was reduced, so which was oxidized?

OpenStudy (aaronq):

which one are you working on?

OpenStudy (anonymous):

g

OpenStudy (anonymous):

I need the answer for g and then: In the equation above, which element was reduced and which element was oxidized?

OpenStudy (aaronq):

The H on the compound is part of the hydroxide ion, which is OH^- since O is -2 and theres a -1 charge on ion, H is +1

OpenStudy (anonymous):

Awesome. so then which was oxidized and which was reduced?

OpenStudy (aaronq):

they're both 0 when not part of a molecule, if one gains e then we say its reduced because it became more negative, the opposite is for oxidized.

OpenStudy (anonymous):

I mean for all of the elements in the equations, it seems like aluminum gained electrons

OpenStudy (anonymous):

!

OpenStudy (aaronq):

we already did everything in the equation